# A Survey of Personalized Large Language Models: Progress and Future Directions

### v1

@@ 7 Conclusions
This survey offers a thorough overview of Personalized Large Language Models (PLLMs), emphasizing personalized responses tailored to individual user data. We introduce a structured taxonomy that categorizes existing approaches into three key technical perspectives: Personalized Prompting (Input Level), Personalized Adaptation (Model Level), and Personalized Alignment (Objective Level), with further subdivisions within each. We also discuss current limitations and propose several promising directions for future research. Our work provides valuable insights and a framework to drive the advancement of PLLMs development.

### v2

@@ 8 Conclusions
This survey offers a comprehensive overview of PLLMs, focusing on personalized responses to individual user data. It presents a taxonomy categorizing approaches into three key perspectives: Personalized Prompting (Input Level), Personalized Adaptation (Model Level), and Personalized Alignment (Objective Level), with further subdivisions. A detailed method summarization is shown in Table 7. We highlight current limitations and suggest future research directions, providing valuable insights to advance PLLM development. Beyond taxonomy, we articulate a vision for memory-centric PLLMs, highlighting three desired capabilities—remember, adapt, and evolve—and analyze their inherent trade-offs among performance, privacy, and efficiency. We further examine domain-specific challenges and discuss future directions. Together, these contributions not only summarize the state of the art but also chart promising pathways for advancing PLLM research and practical deployment.
Appendix A Notation Summary
| Symbol | Description |
| • Spaces | |
| Space of all personalized data | |
| Latent abstraction space (preferences, profiles, tendencies) | |
| Personalization state space (style, policy, reasoning priors, etc.) | |
| Predefined human preference dimensions (e.g., “helpfulness”) | |
| Query space | |
| Output / response space | |
| General or world knowledge available to the LLM | |
| Set of users | |
| • Functions / Mappings | |
| Abstraction / summarization mapping | |
| Personalization encoder (possibly query-aware) | |
| Extraction operator (factual lookup) | |
| Abstraction operator (summary-based answering) | |
| Conditional distribution for generalization queries | |
| • Models | |
| Base large language model (non-personalized) | |
| Personalized model for user | |
| Personalization operator (e.g., prompting, adaptation, alignment) | |
| Tunable parameters (within LLM or added modules, e.g., PEFT/LoRA) | |
| • Operations / Sets | |
| Explicit factual tuples contained in (for Extraction) | |
| • Variables | |
| The -th user | |
| Personalized data of user | |
| Query set of user | |
| The -th query of user | |
| Predicted response for (model output) | |
| Desired personalized output for (ground truth) | |
| Training pair: context/content and reference response | |
| Reference response associated with in training data | |
| Abstracted profile of user | |
| Personalized state for user under query | |